SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the April 23 machine backgrounds video meeting

1. Report from the April 20 BABAR-PEP-II management meeting

    1. SVT and EMC groups confirmed their desired threshold settings, respectively 10 mrad/s and 100mrad/hour. Consensus to begin with such values. For the EMC however, it was not clear whether such a low threshold was realistic, because of noise and temperature drift in the diodes. Tom Mattison indicated that a pin diode will be placed close to the calorimeter but at a lower radius to overcome this problem. Witold made an estimation to compare the proposed 10 mrad/s threshold to what could be expected in the June1999 machine model. (the expectations are typically one order of magnitude higher for ~1 A current with open IR4 collimators)
    2. N. Roe pointed out that with the 10 mrad/s it will be needed to often recalibrate (~ once a shift) the pedestal of the SVT PIN diodes, which is a quick measurement when there is no beam. Instructions will be given so that either the SVT person or the shift liaison makes good use of the occasional no beam periods for this purpose
    3. U. Wienands reported on the rules for decision making and communication during a shift. Detailed guidelines will be written down by Uli, Tom and Iannis about the respective roles of the shift leader and the liaison person.

2. Witold reported on a Montecarlo study to compute the relative contribution of LER and HER in the June 1999 model for the machine. The conclusions are that LER will dominate backgrounds when the IR4 collimators are open (John Seeman indicated that they would be in place by May 7), and when they are closed, HER and LER should contribute about equally. Note that HER contribution is dominated by the the last 20 m of IR2. The HER movable collimator in IR2 should be in in June. Obviously, measuring the impact of IR4 collimators will be a top priority.

3. Artem Kulikov gave a report on the present understanding of pressure variation vs current in HER and LER beams. (work performed with Sam Heifets). In the HER, most pumps behave linearly with current as expected. In the LER, threshold effects are seen, which depend mainly on the bunch current. These thresholds are the clearest in the middle of straight sections 4,8 and 12 (which are really straight) where there is no synchrotron light. Artem indicated that these pumps reading is really due to pressure and not electron currents because of their relatively slow response time to rapid current variation. The rapid rise is interpreted as large desorption induced by electrons , produced by ionisation of the gas by the beam. The process already potentially divergent in itself is further increased by electron secondary emission. The reason these steep thresholds are seen in the middle of the arcs is because of the lack of X rays to desorb: there is lot of gas to get rid of. [In the arcs, the dynamic pressure increase due to the same mechanisms is a factor 3. Note that the dynamic pressure when above the thresholds (typically 1 mA per bunch), is larger in the straights than in the arcs!]

Two ways out: patience (or better, dedicated scrubbing of these zones with high pressure low current operation) and solenoidal field to prevent electrons from travelling across the pipe (note that there is also an ion component to the desorption which will not be affected by B field but which seems to be small). Both will be tried!

  1. Liaison shift schedule. Not yet available. Uli has put on the web a detailed PEP-II schedule for May. See also the startup schedule
  2. Background sensors update. Nothing to report in the absence of Terry and Adrian. Uli mentioned the need for precise operation-oriented documentation for the operators. Tom will provide this for the abort-capable channels.
  3.  

  4. Software week preparation. After last week’s discussion, a detailed agenda for the background oriented session on Friday April 30 is now available. Tom has voluntereed to chair the first and third sessions (background limits and feedback loops), Guy will chair the second on strategies.

7. Due to the software week, there will be no meeting next week, April 30.